OpenStudy (gabby1214):

solve x/7=-7 solution is x=?

OpenStudy (tkhunny):

Have you considered multiplying by 7? Never post a problem without YOUR work showing clearly.

OpenStudy (gabby1214):

Multyplying 7 by 7??

OpenStudy (kendricklamar2014):

Step 1: Simplify both sides of the equation: \[\frac{ 1 }{ 7 } x = -7\]

OpenStudy (kendricklamar2014):

Step 2: Multiply both sides by 7:\[7 * (\frac{ 1 }{ 7 }x) = (7) * (-7)\]

OpenStudy (kendricklamar2014):

@gabby1214 \[7 * -7 = ?\]

OpenStudy (gabby1214):

-49 right?